# PredictaBoard: Benchmarking LLM Score Predictability

### v1

@@ 1 Introduction
This principle extends to frontier AI systems, such as Large Language Models [^1] Many of which can also interpret images and, increasingly, audio and video. (LLMs), which are rapidly being integrated into high-stakes scenarios (Kim et al., 2024; Huang et al., 2024; Javaid et al., 2024). Their growing adoption introduces the potential for high-consequence harm – significant negative impacts on individuals, organisations, or society as a whole– thus requiring robust safeguards to ensure both high performance and safety Hendrycks et al. (2023). Here as well, a key component to risk mitigation is validity predictability, namely, enabling users to anticipate and reject inputs that cause the AI systems to produce undesirable outputs or outcomes Zhou et al. (2024a); Hendrickx et al. (2024); Vafa et al. (2024a); Sec 2 discusses this more in detail.
However, humans struggle to predict when LLMs will be correct (Carlini, 2024), can be biased by prior interactions (Vafa et al., 2024a) and even struggle to evaluate explanations provided by LLMs (Steyvers et al., 2025). Additionally, alignment using Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) and other techniques lead to LLMs potentially deceiving humans more often (Wen et al., 2024; Williams et al., 2024), resulting in unpredictable and unsafe behaviour (Anwar et al., 2024; Zhou et al., 2024b).
@@ 3 Related Work
Humans were found to only marginally beat random guess in predicting GPT-4’s performance (Carlini, 2024). Relatedly, Vafa et al. (2024a) showed humans overestimate LLM future performance based on prior interactions, especially with larger models in high-stakes contexts. They argue that “the best LLM is the one that allows humans to make the most reliable inferences about where it will succeed”, closely aligning with PredictaBoard’s motivation. Zhou et al. (2024b) indicated that human predictions become unreliable as AI systems become more capable and Steyvers et al. (2025) found LLM-produced explanations supporting a statement do not lead humans to reliably assess whether that statement is correct, even when the LLM’s token-level probabilities are calibrated. Finally, in a specific use case, Bansal et al. (2024) reported that the failures of an LLM software engineer agent cannot be reliably anticipated.
@@
Some works trained “assessors” to predict instance-level LLM performance. Kadavath et al. (2022) trained LLMs to predict the probability of succeeding on a question without reference to a specific answer, which performed satisfactorily but struggled with novel tasks. Schellaert et al. (2024) effectively predicts the performance of LLMs on 100+ BIG-bench tasks, outperforming subject systems in confidence, and maintaining predictability across different model sizes, suggesting scalability. Zhou et al. (2022) showed that smaller LLMs can predict performance of larger models on certain tasks, almost halving errors and computational costs. Drapal et al. (2024a) obtained explainable meta-rules from trained assessors to identify regions of predictable performance. Finally, while assessors allow to avoid prompting models on task instances causing failures, their training requires generating a failure/correctness dataset specific to each LLM. Single assessors sharing information across many LLMs Pacchiardi et al. (2024) reduce the instances on which each LLM must be tested. All these techniques can be used as assessors for PredictaBoard.
@@
White-box approaches that rely on model activations also exist (Ferrando et al., 2025). However, in general, the performance of UQ methods is debated (Kapoor et al., 2024); for instance, Pawitan and Holmes (2024) found different methods to extract confidence to be poorly correlated and only partly indicative of correctness. Additionally, in contrast to the anticipative assessors we use as baselines, UQ methods require inputs to be passed through the LLM, making them close to the “integrated rejectors” described in Hendrickx et al. (2024). Nevertheless, PredictaBoard can be used to evaluate these methods too.
@@ 5.2 Out-of-Distribution Evaluation
To assess the robustness of our LLM-assessor pairs, we evaluated them on BBH after training on the train split of MMLU-Pro. Figure 8 replicates Figure 5 for the BBH benchmark. We use the same selection criteria as for the in-distribution results: the union of the top 5 PVR performers at each threshold. The lower values in Figure 8 highlight the difficulty of predicting performance OOD. Additionally, the top pairs differ from the in-distribution ones, suggesting the latter may may not have the highest generalisation power. Other metrics for this OOD scenario are available in App. D.
@@ 6 Conclusion
This paper aimed to release an initial version of the benchmark and allow the community to guide its future extensions, that can possibly include developing assessors that function across multiple LLMs and predicting safety indicators (Zhang et al., 2024; Mazeika et al., 2024), rather than performance. Additionally, comparing with human performance as assessors, as examined in recent studies (Carlini, 2024; Vafa et al., 2024b; Gao et al., 2024; Zhou et al., 2024b), could provide insights into the differences in LLM predictability from automated and human perspectives.
@@ Acknowledgments
7 Limitations
There are some limitations in the current version of PredictaBoard. First, our baselines only include external anticipative assessors. In principle, the PredictaBoard allows for non-external assessors (including self-confidence) and can be easily extended to consider the output of the model as well, using “verifiers” instead of assessors (although not relying on outputs has some advantages, highlighted in §2). Additional conditions could also be considered, such as the invertibility of the assessor—-whether, by using the assessor, one can generate inputs that ensure the model’s performance exceeds a given score, given constraints on the input. This capability would be particularly valuable for red-teaming applications or enhancing explainability (e.g., generating counterfactuals). Considering variations for these additional conditions and others (like the computational cost of the LLM-assessor pair) could allow to study properties of the assessors, such as the exploration of scaling laws for pairs of LLMs and assessors, or explore situations where the assessor cannot be $n$ times more costly than the LLM. We leave all these considerations for future versions and specific competitions.

### v2

@@ 1 Introduction
This principle extends to frontier AI systems, such as Large Language Models [^1] Many of which can also interpret images and, increasingly, audio and video. (LLMs), which are rapidly being integrated into high-stakes scenarios (Kim et al., 2024; Huang et al., 2024; Javaid et al., 2024). Their growing adoption introduces the potential for high-consequence harm – significant negative impacts on individuals, organisations, or society as a whole– thus requiring robust safeguards to ensure both high performance and safety Hendrycks et al. (2023). Here as well, a key component to risk mitigation is validity predictability, namely, enabling users to anticipate and reject inputs that cause the AI systems to produce undesirable outputs or outcomes Zhou et al. (2024a); Hendrickx et al. (2024); Vafa et al. (2024); Sec 2 discusses this more in detail.
However, humans struggle to predict when LLMs will be correct (Carlini, 2024), can be biased by prior interactions (Vafa et al., 2024) and even struggle to evaluate explanations provided by LLMs (Steyvers et al., 2025). Additionally, alignment using Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) and other techniques lead to LLMs potentially deceiving humans more often (Wen et al., 2024; Williams et al., 2024), resulting in unpredictable and unsafe behaviour (Anwar et al., 2024; Zhou et al., 2024b).
@@ 3 Related Work
Humans were found to only marginally beat random guess in predicting GPT-4’s performance (Carlini, 2024). Relatedly, Vafa et al. (2024) showed humans overestimate LLM future performance based on prior interactions, especially with larger models in high-stakes contexts. They argue that “the best LLM is the one that allows humans to make the most reliable inferences about where it will succeed”, closely aligning with PredictaBoard’s motivation. Zhou et al. (2024b) indicated that human predictions become unreliable as AI systems become more capable and Steyvers et al. (2025) found LLM-produced explanations supporting a statement do not lead humans to reliably assess whether that statement is correct, even when the LLM’s token-level probabilities are calibrated. Finally, in a specific use case, Bansal et al. (2024) reported that the failures of an LLM software engineer agent cannot be reliably anticipated.
@@
Some works trained “assessors” to predict instance-level LLM performance. Kadavath et al. (2022) trained LLMs to predict the probability of succeeding on a question without reference to a specific answer, which performed satisfactorily but struggled with novel tasks. Schellaert et al. (2024) effectively predicts the performance of LLMs on 100+ BIG-bench tasks, outperforming subject systems in confidence, and maintaining predictability across different model sizes, suggesting scalability. Zhou et al. (2022) showed that smaller LLMs can predict performance of larger models on certain tasks, almost halving errors and computational costs. Drapal et al. (2024a) obtained explainable meta-rules from trained assessors to identify regions of predictable performance. Finally, while assessors allow to avoid prompting models on task instances causing failures, their training requires generating a failure/correctness dataset specific to each LLM. Single assessors sharing information across many LLMs (Pacchiardi et al., 2024) reduce the instances on which each LLM must be tested. All these techniques can be used as assessors for PredictaBoard.
Factuality detection and correctness prediction with model internals
Model internal activations have been used to detect statement truthfulness (Burns et al., 2022; Azaria and Mitchell, 2023; Marks and Tegmark, 2023; Burger et al., 2024), deception (Goldowsky-Dill et al., 2025) and hallucination (Ferrando et al., 2025). However, these works considered entire statements (or question–answer pairs), making them inherently non-anticipative. A few studies (Kadavath et al., 2022; Ferrando et al., 2025) instead rely solely on the internal activations produced by the question to predict the correctness of the answer the model will generate; these can be viewed as assessors based on model internals, and can be evaluated within the PredictaBoard framework, even though the baseline approaches we present here use only model-independent features of the input prompt. While tapping into the internal embedding may boost predictability, it hinges on (and thus exposes) the model’s own capacity to form confidence measures internally. In contrast, external assessors identify and exploit model-independent features that capture the intrinsic demands of the task.
@@
However, in general, the performance of UQ methods is debated (Kapoor et al., 2024); for instance, Pawitan and Holmes (2024) found different methods to extract confidence to be poorly correlated and only partly indicative of correctness. Additionally, in contrast to the anticipative assessors we use as baselines, UQ methods require inputs to be passed through the LLM, making them close to the “integrated rejectors” described in Hendrickx et al. (2024). Nevertheless, PredictaBoard can be used to evaluate these methods too.
Reward models
Reward modelling typically evaluates a model’s output against “soft” criteria such as toxicity (Faal et al., 2023) or alignment with user intent (Ziegler et al., 2019; Ouyang et al., 2022); instead, the assessors subject of PredictaBoard predict the objective correctness of an answer, even though the validity notion PredictaBoard uses could be extended to include toxicity or related criteria. Additionally, PredictaBoard considers predicting the score from the question only, while reward modelling relies on the generated answer and is therefore more closely related to score modelling (automated grading, LLM-as-a-judge, Zhu et al., 2023). Moreover, reward models are learnt independently of a particular LLM, while PredictaBoard’s assessors are specialized to anticipate a specific LLM’s behaviour.
@@ 5.2 Out-of-Distribution Evaluation
To assess the robustness of our LLM-assessor pairs, we evaluated them on BBH after training on the train split of MMLU-Pro. Figure 8 replicates Figure 5 for the BBH benchmark. We use the same selection criteria as for the in-distribution results: the union of the top 5 PVR performers at each threshold. The lower values in Figure 8 highlight the difficulty of predicting performance OOD. Additionally, the top pairs differ from the in-distribution ones, suggesting the latter may not have the highest generalisation power. Other metrics for this OOD scenario are available in App. D.
@@ 6 Conclusion
This paper aimed to release an initial version of the benchmark and allow the community to guide its future extensions, that can possibly include developing assessors that function across multiple LLMs and predicting safety indicators (Zhang et al., 2024; Mazeika et al., 2024), rather than performance. Additionally, comparing with human performance as assessors, as examined in recent studies (Carlini, 2024; Vafa et al., 2024; Gao et al., 2024; Zhou et al., 2024b), could provide insights into the differences in LLM predictability from automated and human perspectives.
7 Limitations
There are some limitations in the current version of PredictaBoard. First, our baselines only include external anticipative assessors. In principle, PredictaBoard allows for non-external assessors (including self-confidence and latent LLM’s embeddings, as discussed in Sec. 3) and can be easily extended to consider the output of the model as well, using “verifiers” instead of assessors (although not relying on outputs has some advantages, highlighted in Sec. 2). Additional conditions could also be considered, such as the invertibility of the assessor—-whether, by using the assessor, one can generate inputs that ensure the model’s performance exceeds a given score, given constraints on the input. This capability would be particularly valuable for red-teaming applications or enhancing explainability (e.g., generating counterfactuals). Considering variations for these additional conditions and others (like the computational cost of the LLM-assessor pair) could allow to study properties of the assessors, such as the exploration of scaling laws for pairs of LLMs and assessors, or explore situations where the assessor cannot be $n$ times more costly than the LLM. We leave all these considerations for future versions and specific competitions.